OpenStudy (anonymous):

Mary starts from her house, walks 80 meters south, and stops to chat with her aunt on the sidewalk. After chatting for a few minutes, she walks 125 meters to the west to meet a friend and then walks 45 meters north with her friend to a cafe for lunch. Overall, she takes 10 minutes to do all this. What is Mary’s average speed?Select one of the options below as your answer:A. 8 meters/secondB. 1.25 meters/secondC. 25 meters/secondD. 4.5 meters/secondE. 0.42 meters/second

OpenStudy (anonymous):

Distance traveled = 80 + 125 + 45 = 250 meters Time = 10 min x 60 s/min = 600 seconds Speed = Distance / Time = 250 / 600 = 0.42 meters/second or E

OpenStudy (anonymous):

We cannot answer without proper information.We can answer only if you say how much time did Mary chat with her aunt.